1. 2.08 poimts ou invest S4,C0 far three yeere al 9 pencent a. whal ะ 1ne vaue uf sour invest mert alter cne year? Muhply S4UU-1.uy. b. ttha valuayar rwaslmantoyaMubphy yoursarn pat a .. (Round your anawer to 2 decimal placea.) G. What s the vaiue uf your investment after three yeers? Mutipy yuranswer to pert s by 1. This gives ycur fin anwer. Round your answer to 2 decimal places.l

please answer all questions!!
Answer #1

Initial Investment amount = $ 4,000

Interest rate receivable on investment = 9%

a. Value of Investment after one year = Initial investment * (1 + interest) = 4,000 * 1.09 = $ 4,360

b. Value of Investment after two year = Initial investment *(1+interest of 1st year) *(1+interest of 2nd year)

= 4,000 *(1.09) *(1.09)

= $ 4,752.40

c. Value of Investment after three year = Initial investment *(1+interest of 1st year) *(1+interest of 2nd year) *(1+interest of 3rd year)

= 4,000 *(1.09) *(1.09) *(1.09)

= $ 5,180.12